OpenStudy (mcm128):

Can someone please show me how to do this? Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.) μ = 104; σ = 16 P(x ≥ 120) =

OpenStudy (mathmate):

First find the Z-score using \(Z=(x-\mu)/\sigma\)=(x-104)/16 =(120-104)/16=1 [ for x=120] P(x\(\ge 120\))=P(Z>1) Then look up the probability distribution (table or calculator) to find the probability of Z > 1.00 See probability distribution table below: http://www.stat.purdue.edu/~mccabe/ips4tab/bmtables.pdf Look for the row where z=1.0. The first column is z=1.00 (what you need), the second column is z=1.01, etc. The table entry represent probabilities.